OpenStudy (anonymous):

Any help would be awesome!(: The dielectric constant of ruby mica is 5.4. What would be the speed of light as it passes through ruby mica?

OpenStudy (schrodingers_cat):

The Equation your looking for is v = c/n n = refractive index c= speed of light k = dielectric constant Use the relation k = n^2 So, if you plug in the numbers you get, v = (3 x 10^8)/(sqrt(5.4)) = 129.09 x10^6 m/s I hope this helps :)

OpenStudy (anonymous):

wow.. schrodingers cat.. are u dead or alive? :D

OpenStudy (anonymous):

It did help but on the problem, I have answer choices and neither match up /: there's 3.2x10^3 m/s, 9.4x10^4 m/s, 5.6x10^7 m/s, and 1.3x10^8 m/s. I'm thinking it should be the last one? @Schrodingers_Cat

OpenStudy (schrodingers_cat):

Yes. I would try the last one I you round up you should get roughly 1.3 x 10^8 m/s Hope this helps :)

OpenStudy (anonymous):

Okay, great. Thank you so much! (: